e following C++ program and provide the correct output: int x = 20, count=0; do { cout << x; x = 4* (x + 5)/5; count++; }whil trace
Answer #1

Answer:

output: 2020

it prints 20 value 2 times.it is not 2020

tracing:

x is an integer and having value 20

x=20

count=0 and it is also a integer

There a do-while loop.

Now loop is initiated.

And loop iterates for two times because while (count<2).

1st iteration

cout<<x

It prints 20.

x=4*(x+5)/5

x=4*(20+5)/5

x=4*25/5

x=100/5=20

So again x=20

In next line count incremented so count =1

count<2 --> 1<2,true

2nd iteration

cout<<x

It prints 20 again.

Again x=4*(x+5)/5

x=20

count incremented so count=2

count<2 ,2<2 false so

The final output is 2020.